Island Novelties, Inc., of Palau makes two products—Hawaiian Fantasy and Tahitian Joy. Each products selling price, variableReq 1A Reg 1B Reg 2A Reg 2B Assuming the sales mix given above, do the following: Prepare a contribution format income statemReg 1A Reg 1B Req 2A Req 2B Assuming the sales mix given above, do the following: Compute the companys break-even point in dReq 1A Reg 1B Reg 2A Reg 2B The company has developed a new product called Samoan Delight that sells for $50 each and that haReg 1A Reg 1B Reg 2A Req 2B The company has developed a new product called Samoan Delight that sells for $50 each and that ha
Answer #1
1-a.
Island Novelties, Inc.
Contribution Income Statement
Hawaiian Fantasy Tahitian Joy Total
Amount % Amount % Amount %
Sales 680000 100% 720000 100% 1400000 100%
Variable expenses 442000 65% 216000 30% 658000 47%
Contribution margin 238000 35% 504000 70% 742000 53%
Fixed expenses 651900
Net operating income 90100
1-b.
Break-even point in dollars = Fixed expenses / Contribution margin ratio of company = 651900 / 53% = 1230000
Margin of safety in dollars = Sales - Break even point in dollars = 1400000 - 1230000 = 170000
Margin of safety percentage = Margin of safety in dollars / Sales = 170000 / 1400000 = 12.1%
2-a.
Island Novelties, Inc.
Contribution Income Statement
Hawaiian Fantasy Tahitian Joy Samoan Delight Total
Amount % Amount % Amount % Amount %
Sales 680000 100.0% 720000 100.0% 600000 100.0% 2000000 100.0%
Variable expenses 442000 65.0% 216000 30.0% 360000 60.0% 1018000 50.9%
Contribution margin 238000 35.0% 504000 70.0% 240000 40.0% 982000 49.1%
Fixed expenses 651900
Net operating income 330100
2-b.
Break-even point in dollars = Fixed expenses / Contribution margin ratio of company = 651900 / 49.1% = 1327699
Margin of safety in dollars = Sales - Break even point in dollars = 2000000 - 1327699 = 672301
Margin of safety percentage = Margin of safety in dollars / Sales = 672301 / 2000000 = 33.6%